# Collision-Free Robot Navigation in Crowded Environments using Learning based Convex Model Predictive Control

### v2

@@ I INTRODUCTION
  • We have made our work open source by publishing both the implementation and benchmark data in a code repository accessible at https://github.com/sunnyhello369/flat_ped_sim.git.
@@ III-C Action Space Design Based on Convex Region
When designing the action space, we first ensure that the sampling of reference points is confined within the convex region. Then, centering on the robot’s current position $O$ and considering its kinematic limits, we define two circular areas. These circles, intersected with the convex region, represent our short-term and long-term action spaces, as illustrated by the red areas in Fig. 2. The policy network is designed to output two sets of two-dimensional data. These data, after sigmoid activation, yield $\hat{a}_{t}=\left\{\left(\alpha_{t}^{s},\beta_{t}^{s}\right),\left(\alpha_{t}^{l},\beta_{t}^{l}\right)|\alpha_{t},\beta_{t}\in\left(0,1\right)\right\}$.
@@ III-D Model Predictive Control Formulation
Using the method mentioned in Section III-C, we get the convex region from point cloud data, with its vertices $P_{j}^{c},j=\left\{1,...,rnum_{v}\right\}$ arranged in clockwise direction. This region constrains the MPC-optimized points to enhance safety by reducing the risk of collisions with obstacles. This constraint is efficiently met by determining whether the trajectory point $Q_{i}$ is inside the convex region through the vector cross product method $\overrightarrow{Q_{i}P_{j}^{c}}\times\overrightarrow{Q_{i}P_{j+1}^{c}}\leq 0$.
During navigation, this vector method is also used to verify if the goal endpoint lies within the convex region. If so, this endpoint is directly used as the long-term reference point $Q_{t}^{l}$ , which promotes early training by facilitating the acquisition of high-reward trajectories. Additionally, we introduce a final state stop cost $\mathrm{cost}_{stop}$, defined in Equation 3, into the MPC’s cost function when the goal is within the convex region. This cost aims to minimize both velocity $V_{N}=\left(v_{x}^{N},v_{y}^{N}\right)$ and acceleration $A_{N}=\left(a_{x}^{N},a_{y}^{N}\right)$ at the destination, ensuring a controlled and safe termination of movement.
The short-term and long-term reference points are used to guide the positions of the first and last control period states predicted by the MPC. The cost function incorporates discrepancies between MPC-predicted points and reference points to ensure fidelity to the reference trajectory. Ultimately, the MPC is defined as a quadratic programming problem, formulated in Equation 4, optimizing the control sequence $u_{0:N-1}^{*}$. The optimization process generates a sequence of MPC-optimized points $Q_{i}^{*},i=\left\{1,...,N\right\}$, forming the MPC-optimized trajectory, as depicted in Fig. 4.
$\mathrm{cost}_{stop}=\begin{cases}\begin{aligned} &w_{vend}\|V_{N}\|^{2}\\ &+w_{aend}\|A_{N}\|^{2}\end{aligned}&\text{if goal in convex}\\ 0&\text{otherwise}\end{cases}$ (3)
@@
$\quad+w_{smooth}\sum_{k=0}^{N-1}{\left\|u_{k}\right\|^{2}}+\mathrm{cost}_{stop}$
s.t. $x_{0}=x_{init},$
@@
$\forall i\in\{1,...,N\};\forall j=\left\{1,...,rnum_{v}-1\right\}.$
@@ III-F Reward Function Formulation
The collision penalty $r_{t}^{o}$, derived from distance data $scan$ collected by LiDAR, is dynamically adjusted. The penalty decreases exponentially as the distance increases, controlled by a positive decay factor $w_{obs}$.
@@
To mitigate sparse reward, we introduce the reward $r_{t}^{a2}$ , based on the Euclidean distance $d_{t}$ between the robot and its destination.
$r_{t}^{a2}=\begin{cases}0&\text{if }t=1\\ w_{approach2}(d_{t}-d_{t-1})&\text{otherwise}\end{cases}$ (8)
Short-term and long-term reference points, $Q_{t}^{s}$ and $Q_{t}^{l}$, discussed in Section III-C, might not be practical to reach. However, the MPC-optimized positions $Q_{i}^{*}$ , outlined in Section III-D, comply with kinematic constraints and provide more feasible references. Thus, to penalize the discrepancy between reference points and MPC-optimized points, we introduce $r_{t}^{f}$.This reward encourages the generation of reachable reference points.
@@
Significant changes in reference points across frames can result in deviations between MPC-optimized trajectories at consecutive steps. Consequently, the agent requires increased control efforts to mitigate the influence of previous frames. To address this, the penalty $r_{t}^{c}$ is introduced, which can also facilitate rapid learning of smooth direction changes during the initial training phase.
@@ III-G Network Architecture
Figure 5: Network Architecture: The value network $V_{\phi}$ and policy network $\pi_{\theta}$ are structured as four-layer fully connected networks. The policy network is augmented with a logarithmic standard deviation parameter ($\ln\sigma_{t}$) for each action dimension. This configuration primarily facilitates the generation of stochastic policies by allowing actions to be sampled from a Gaussian distribution, where $\sigma_{t}$, derived from $\ln\sigma_{t}$, modulates the balance between exploration and exploitation.
@@ IV-A2 Implementation Details
Our study utilizes the Arena-Rosnav framework [29], developed within the ROS architecture, to enable reliable robotic navigation solutions. We also adopt the implementation of Proximal Policy Optimization (PPO) [30] from Elegant-RL [31], which provides tools for network customization and training process monitoring. The simulation experiments are conducted on a high-performance PC equipped with an Intel Core i7-7700K CPU, an Nvidia RTX 2080Ti GPU, and 32GB of RAM.
@@ IV-D Evaluation
Following Stages 5-7’s training in tight, dynamic crowded environments, both our method and TEB method are evaluated in 1,000 test scenarios for each Stage, with results detailed in Table IV.
@@
Figure 6: Comparative Navigation Process: Our Method and TEB in Selected Test Scenarios from Stages 5, 6, 7. The three images, arranged from top to bottom, respectively depict the robot’s navigation process during the test scenarios of Stages 5, 6, and 7. Light blue polygons for static obstacles, dark blue line marking robot trajectory, a green star for the goal, and cyan numbers indicate the time used by robot to move to this position, in seconds. Red lines with adjacent black numbers represent dynamic obstacles’ paths and timing move to this position, respectively.
@@ IV-E Qualitative Analysis
Figure 7 illustrates the navigation process, reflecting the iterative trajectory optimization procedure depicted in Figure 3. The robot updates the convex obstacle-free region with the latest LiDAR data. Subsequently, long-term and short-term reference points are sampled within this convex region to form the reference trajectory. The MPC then tracks this reference trajectory to solve the local navigation problem, until the goal endpoint is within the latest convex region and is directly adopted as the long-term reference point. The process aims to promote safety by constraining both the reference trajectory and robot motion within the overlapping convex regions. Additionally, the frequent updating of reference points significantly enhances dynamic safety, ensuring a robust response to environmental changes.
@@ V CONCLUSIONS
This paper introduces a DRL-based navigation framework, which consists of the design of action and state spaces, reward function and network architecture. In terms of action space design, our method integrates lidar-generated convex obstacle-free region to formulate the action space that includes both short-term and long-term reference points. For the reward function, we devise a composite reward function enriched with intermediate rewards. We conduct training and experiments across seven Stages. The experimental results conclusively demonstrate that the proposed method significantly enhances robotic performance in both static and dynamic environments, notably excelling in dynamic and crowded environments.

### v3

@@ I INTRODUCTION
  • We have made our work open source by publishing both the implementation and benchmark data111 https://github.com/sunnyhello369/flat_ped_sim.git.
@@ III-C Action Space Design Based on Convex Region
When designing the action space, we first ensure that the sampling of reference points is confined within the convex region. Then, centering on the robot’s current position $O$ and considering its kinematic limits, we define two circular areas. These circles, intersected with the convex region, represent our short-term and long-term action spaces, as illustrated by Fig. 2. The policy network is designed to output two sets of two-dimensional data. These data, after sigmoid activation, yield $\hat{a}_{t}=\left\{\left(\alpha_{t}^{s},\beta_{t}^{s}\right),\left(\alpha_{t}^{l},\beta_{t}^{l}\right)|\alpha_{t},\beta_{t}\in\left(0,1\right)\right\}$.
@@ III-D Model Predictive Control Formulation
Using the method mentioned in Section III-C, we get the convex region from point cloud data, with its vertices $P_{j}^{c},j=\left\{1,...,rnum_{v}\right\}$ arranged in clockwise direction. $rnum_{v}$ is a hyperparameter that describes the fixed number of vertices and will be introduced at Section III-E. This region constrains the MPC-optimized points to enhance safety by reducing the risk of collisions with obstacles. This constraint is efficiently met by determining whether the trajectory point $Q_{i}$ is inside the convex region through the vector cross product method $\overrightarrow{Q_{i}P_{j}^{c}}\times\overrightarrow{Q_{i}P_{j+1}^{c}}\leq 0$.
During navigation, this vector method is also used to verify if the goal endpoint lies within the convex region. If so, this endpoint is directly used as the long-term reference point $Q_{t}^{l}$ , which promotes early training by facilitating the acquisition of high-reward trajectories. Additionally, we introduce a final state stop cost $\mathrm{cost}_{stop}$, into the MPC’s cost function when the goal is within the convex region. This cost aims to minimize both velocity $V_{N}=\left(v_{x}^{N},v_{y}^{N}\right)$ and acceleration $A_{N}=\left(a_{x}^{N},a_{y}^{N}\right)$ at the destination, ensuring a controlled and safe termination of movement.
The short-term and long-term reference points are used to guide the positions of the first and last control period states predicted by the MPC. The cost function incorporates discrepancies between MPC-predicted points and reference points to ensure fidelity to the reference trajectory. Ultimately, the MPC is defined as a quadratic programming problem, formulated in Equation 3, optimizing the control sequence $u_{0:N-1}^{*}$. $w_{track}$ and $w_{smooth}$ are the weights of the tracking error term and the smoothing term in the cost function respectively. $w_{vend}$ and $w_{aend}$ are the final state velocity and acceleration weights of the optimised $\mathrm{cost}_{stop}$. The optimization process generates a sequence of MPC-optimized points $Q_{i}^{*},i=\left\{1,...,N\right\}$, as depicted in Fig. 4.
@@
$+w_{smooth}\sum_{k=0}^{N-1}{\left\|u_{k}\right\|^{2}}+\cos t_{stop}$
$\mathrm{cost}_{stop}$ =$=\begin{cases}\begin{split}&w_{vend}\|V_{N}\|^{2}\\ &+w_{aend}\|A_{N}\|^{2}\end{split}&\text{if goal in convex}\\ 0&\text{otherwise}\end{cases}$
$\mathrm{s}.\mathrm{t}.$ $x_{0}=x_{init}$
@@
$\forall i\in\{1,...,N\};\forall j=\left\{1,...,rnum_{v}-1\right\}$
@@ III-F Reward Function Formulation
The collision penalty $r_{t}^{o}$, derived from distance data $scan$ collected by LiDAR, is dynamically adjusted. The penalty decreases exponentially as the distance increases. This penalty is controlled by a positive decay factor $w_{obs}$ and two negative terms $r_{obs}$ and $r_{collision}$.
@@
To mitigate sparse reward, we introduce the reward $r_{t}^{a}$ , based on the Euclidean distance $d_{t}$ between the robot and its destination which controlled by a positive factor $w_{approach}$.
$r_{t}^{a}=\begin{cases}0&\text{if }t=1\\ w_{approach}(d_{t}-d_{t-1})&\text{otherwise}\end{cases}$ (7)
We also tried adding guide reward based on global path to $r_{t}^{a}$.
Short-term and long-term reference points, $Q_{t}^{s}$ and $Q_{t}^{l}$, discussed in Section III-C, might not be practical to reach. However, the MPC-optimized positions $Q_{i}^{*}$ , outlined in Section III-D, comply with kinematic constraints and provide more feasible references. Thus, to penalize the discrepancy between reference points and MPC-optimized points, we introduce $r_{t}^{f}$.This reward encourages the generation of reachable reference points.This penalty is controlled by two negative penalty factor $w_{feasible}^{s}$ and $w_{feasible}^{l}$.
@@
Significant changes in reference points across frames can result in deviations between MPC-optimized trajectories at consecutive steps. Consequently, the agent requires increased control efforts to mitigate the influence of previous frames. To address this, the penalty $r_{t}^{c}$ is introduced, which can also facilitate rapid learning of smooth direction changes during the initial training phase.This penalty is controlled by two negative penalty factor $w_{change}^{s}$ and $w_{change}^{l}$.
@@ III-G Network Architecture
Figure 5: Network Architecture: The value network $V_{\phi}$ and policy network $\pi_{\theta}$ are structured as four-layer fully connected networks. The policy network is augmented with a logarithmic standard deviation parameter ($\ln\sigma_{t}$) for each action dimension. This configuration primarily facilitates the generation of stochastic policies by allowing actions to be sampled from a Gaussian distribution.
@@ IV-A2 Implementation Details
Our study utilizes the Arena-Rosnav framework [29], developed within the ROS architecture, to enable reliable robotic navigation solutions. We also adopt the implementation of Proximal Policy Optimization (PPO) [30] from Elegant-RL [31], which provides tools for network customization and training process monitoring. The simulation experiments are conducted on a high-performance PC equipped with an Intel Core i7-770to0K CPU, an Nvidia RTX 2080Ti GPU, and 32GB of RAM.In addition, we tried multi-GPU training experiments on a computer with an i7-6800k CPU and two Nvidia RTX 1080 Ti GPUs.
@@ IV-D Evaluation
Following Stages 5-7’s training in tight, dynamic crowded environments, both our method and TEB method are evaluated in 1,000 test scenarios for each Stage, with results detailed in Table IV.In order to use the same kinematic model as our method, we debug and experiment on the basis of omni-directional TEB [^2] https://github.com/pingplug/teb_local_planner/tree/omni_type.
@@
Figure 6: Comparative Navigation Process: Our Method and TEB in Selected Test Scenarios from Stages 5, 6, 7. The three images, arranged from top to bottom, respectively depict the robot’s navigation process during the test scenarios of Stages 5, 6, and 7. Dark blue line marking robot trajectory, a green star for the goal, and cyan numbers indicate the time used by robot to move to this position, in seconds. Red lines with adjacent black numbers represent dynamic obstacles’ paths and timing move to this position, respectively.
@@ IV-E Qualitative Analysis
Figure 7 illustrates the navigation process, reflecting the iterative trajectory optimization procedure depicted in Figure 3. The robot updates the convex obstacle-free region with the latest LiDAR data. Subsequently, long-term and short-term reference points are sampled within this convex region to form the reference trajectory. The MPC then tracks this reference trajectory to solve the local navigation problem, until the goal endpoint is within the latest convex region and is directly adopted as the long-term reference point. The process aims to promote safety by constraining both the reference trajectory and robot motion within the overlapping convex regions.
@@ V CONCLUSIONS
This paper introduces a DRL-based navigation framework, which consists of the design of action and state spaces, reward function and network architecture. In terms of action space design, our method integrates lidar-generated convex obstacle-free region to formulate the action space that includes both short-term and long-term reference points. For the reward function, we devise a composite reward function enriched with intermediate rewards. The experimental results conclusively demonstrate that the proposed method significantly enhances robotic performance in both static and dynamic environments, notably excelling in dynamic and crowded environments.